# CrEIMBO: Cross Ensemble Interactions in Multi-view Brain Observations

### v1

@@ 3 Problem Formulation and Approach
Fitting this model requires identifying the global operators $\{\bm{f}_{k}\}_{k=1}^{K}$ and the functional ensembles $\{\bm{A}^{d}\}_{d=1}^{D}$. For this, we employ an alternating approach, where we iterate between updating the global operators, $\{\bm{f}_{k}\}_{k=1}^{K}$, and infer, for each session $d$, 1) the ensemble compositions $\bm{a}:=\bm{A}^{d}$, 2) their dynamics $\bm{x}$, and 3) the sub-circuits’ temporal coefficients $\bm{c}_{t}$. These steps are iterated until the parameters converge.
@@
Ensemble update: We update $\bm{a}$ per-row (neuron) $n$ ($\bm{a}_{n}$). Following the work of [17], we infer the sparse ensemble structures with graph-driven re-weighted $\ell_{1}$ regularization that groups together neurons with shared activity patterns, while pushing apart neurons who do not behave similarly. The graph for each area $j$, $\bm{h}^{j}\in\mathbb{R}^{n_{j}\times n_{j}}$, is calculated based on a data-driven Gaussian kernel that measures the temporal similarity between neurons, such that $\bm{h}^{j}_{n1,n2}=exp(\frac{\|\bm{y}^{j}_{n1}-\bm{y}^{j}_{n2}\|_{2}^{2}}{\sigma_{h}})$ where $\sigma_{h}$ is a hyperparameter. Specifically, for each row $\bm{a}_{n,:}$ and considering only the ensembles in region $j_{n}$ that neuron $n$ is part of,
$\widehat{\bm{a}}_{n,:}=\arg\min_{\bm{a}_{n:}}\|\bm{y}_{n:}-\bm{a}_{n:}\bm{x}_{j_{n}:}\|_{2}^{2}+\sum_{j=1}^{p}\bm{\lambda}^{d}_{n,j}|\bm{a}_{n,j}|\quad\textrm{ where }\quad\bm{\lambda}_{n,j}=\frac{\beta_{1}}{\beta_{2}+|\widehat{\bm{a}}_{n,j}|+\beta_{3}|\bm{h}_{n:}^{j}\widehat{\bm{a}}_{:j}|}.$
@@
Here, $\bm{x}_{j_{n},:}$ is the activity of the ensembles of the $j_{n}^{th}$ region and $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are scalars that control the effect of the graph on the regularization. The graph adjacency $\bm{h}_{n:}\widehat{\bm{a}}_{:j}$ weights the sparse membership in the $j$-th ensemble based on temporal similarity. Specifically, if $\bm{h}_{n:}\widehat{\bm{a}}_{:j}$ is high (i.e., the neighbors neuron $n$ on the graph are in the current estimate of the $j$-th ensemble), $\bm{\lambda}_{n,j}$ is small, resulting in lower sparsity regularization on neuron $n$ for that ensemble, promoting it to be included. Alternatively, if this correlation is small, the regularization weight $\bm{\lambda}_{n,j}$ is large, and neuron $n$ will be more likely not to be included in ensemble $j$.
@@
${\widehat{\bm{x}}_{t},\widehat{\bm{c}}_{t}=\arg\min_{\bm{x}_{t},\bm{c}_{t}}\|\bm{y}_{t}-\bm{a}\bm{x}_{t}\|_{2}^{2}+\left\|\bm{x}_{t}-\left(\sum_{k=1}^{K}\bm{f}_{k}\bm{c}_{k,(t-1)}\right)\bm{x}_{t-1}\right\|_{2}^{2}+\lambda_{x}\|\bm{x}_{t}\|_{1}+\lambda_{c}\|\bm{c}_{t}\|_{1}}$ (1)
where $\lambda_{c}\in\mathbb{R}$ and $\lambda_{x}\in\mathbb{R}$ are regularization weights.
@@
$\bm{F}^{\textrm{all}}_{n}=\arg\min_{{\bm{F}^{\textrm{all}}}_{j}}\|\bm{y}_{n}-\bm{F}_{n}^{\textrm{all}}\bm{cx}\|+\lambda_{F}\|\bm{F}^{\textrm{all}}_{j}\|_{1}+\sum_{{k_{1}},{k_{2}},({k_{1}}\neq{k_{2}})}\rho({\bm{f}}_{k_{1}},{\bm{f}}_{k_{2}})$ (2)
@@
where $\bm{F}^{\textrm{all}}$ is an horizontal concatenation of all $\bm{f}$s, and $\sum_{{k_{1}},{k_{2}}}\rho(\bm{f}_{k_{1}},\bm{f}_{k_{2}})$ is a de-correlation term used to ensure that distinct $\bm{f}$s are not too similar, Above, $\bm{cx}_{t}:=(\bm{c}^{batch}_{t}\otimes[1]_{1\times p})\circ([1]_{1\times K}\otimes\bm{x}^{batch}_{t})$, and $\lambda_{F}$ is the weight on the sparsity-regularization on $\bm{F}$ and “batch” means that this operation is in practice taken on a random subset of sessions for computational complexity considerations.
Figure 2: A-E: Testing CrEIMBO over 312 iterations of random initializations and parameters (see Sec. D) reveals high correlations between ground truth and fit components for all unknowns, $\bm{Y}$, $\bm{A}$, $\bm{x}$, $\bm{F}$, $\bm{c}$ (subplots A-E respectively). F-G: A comparison between the ensemble trajectory ($\bm{X}$) and the ensemble compositions ($\bm{A}$) as identified by CrEIMBO and other approaches compared to the ground-truth. Details about the other approaches are found in Section F. Additional sessions in Figure 5. H-J: $-\log(MSE)$ between the ground truth latent dynamics, ensemble composition, and sub-circuit, to the corresponding components as identified by CrEIMBO vs. other approaches. The sub-circuits were re-ordered for each method to maximize correlation with the ground-truth circuits using Scipy’s linear_sum_assignment [4].
@@
Figure 3: CrEIMBO identifies cross-regional neural sub-circuits underlying multi-session human brain recordings. A: Two exemplary observations ($\bm{Y}$, right) compared to their reconstruction by crEIMBO ($\widehat{\bm{Y}}$, left). B: Reconstruction performance in terms of $R^{2}$ and relative error across all sessions. C: The sub-circuit coefficients for 3 exemplary subjects. D: The standard deviation (std) of the coefficients over time across the two exemplary sessions. E: Examplary identified sparse ensemble matrices ($\{\bm{A}^{d}\}_{d=1}^{D}$) for the observations from A. F: Relative std of the coefficients across four subjects reveals usage of similar sub-circuits. G: The identified ensembles’ activity ($\bm{X}^{d}$) for the two exemplary sessions. H: The identified sub-circuits. Width indicates effect magnitude. Each circle corresponds to an ensemble with its color indicating the ensemble’s area. Black arrows near $\bm{f}_{8}$ and $\bm{f}_{11}$ highlight within-region interactions. I: Pairwise correlations between within-session sub-circuit coefficients (left, middle) and average within-session correlations across all sessions with std marked by the size of a “+” marker. J: Average and std of connection values in $f$s. K: Number of non-zero connections per sub-circuit $\bm{f}_{k}$.
@@ 4 Experiments
4.1 CrEIMBO recovers Ground Truth Components in Synthetic Data
We first assess CrEIMBO’s ability to recover ground-truth components from synthetic data. We generated $K=3$ sub-circuits represented by a set of 3 rotational matrices (Fig. 4B). These sub-circuits capture interactions among $p=3$ ensembles from $J=3$ regions ($p_{j}=1$ for all $j=1,\dots,J$). Each region consists of a random number of neurons drawn from a uniform distribution between $4$ to $9$. We finally generated the data by simulating 5 sessions with different neuron identities and varying numbers of neurons per region(Fig. 4A and D).
@@ 4.1 CrEIMBO recovers Ground Truth Components in Synthetic Data
CrEIMBO was able to recover the ground truth components with high accuracy, as measured by correlation with the ground truth, across $312$ random noisy initialization repeats with random parameters from a range described in 1. Accurate recover included all parameters: the reconstructed observations $\bm{y}$ (Fig. 2A), ensemble composition $\bm{A}$ (Fig. 2B) and activity $\bm{x}$ (Fig. 2C), sub-circuits $\bm{F}$ (Fig. 2D), and coefficients $\bm{c}$ (Fig. 2E). Comparing CrEIMBO’s results to components identified by other methods (baselines detailed in F), we found that approaches that either separate the ensembles and dynamics identifications steps, analyze each session separately, or do not consider the localized structure of the ensembles, fail to recover the ground-truth components, i.e., the resulting fit model parameters are less correlated with the ground-truth observations (Fig. 2F-J, 5).
@@
Next, we tested CrEIMBO on a richer synthetic example with $D=15$ sessions. Each session had a varying number of neurons (a random integer between $14$ and $19$ per region from a uniform distribution) and distinct sub-circuit coefficients (see Fig. 8 for the generated ground truth components and 6 for regional distributions and masking). The sub-circuits were set as rotational matrices, similar to the previous experiment. Over $204$ random initializations with different parameters (see Sec. D) we again found that CrEIMBO robustly identifies the ground truth latent dynamics, ensemble compositions, and sub-circuits. Moreover, CrEIMBO accurately reconstructed the observations with high correlation to the ground truth across all initializations (Fig. 8E) with low relative error (Fig. 8F).
We further tested CrEIMBO on more advanced data that included $D=40$ sessions with a maximum of $4$ regions and $p_{j}=3$ ensembles per region (Fig. 9). We found that CrEIMBO manages to recover the components with high correlation with the ground-truth (Fig 10).
4.2 CrEIMBO Discovers Multi-Regional Dynamics in Human Electrophysiology Data
We then tested CrEIMBO on human neural recordings from a high-density electrode array provided by [12]. The data consists of neural activity from $p=10$ brain areas in the human medial temporal and medial frontal lobes, recorded while subjects performed a screening task (detailed in [12]). The data encompasses $21$ subjects across $D=41$ non-simultaneous sessions. For each subject (except Subject 19), the data offers two recording sessions: 1) a “Screening” session to identify natural images that elicit strong neural responses for that subject and 2) a “Sternberg test” session to assess working memory using the images identified in the screening session. For Subject 19, the data includes only the Sternberg session. The data encompasses multiple regions with limited cross-session regional overlap, a few missing samples per session, and variability in the distribution and count of neurons over areas, with some areas having no neurons in some trials or subjects (Fig. 11, 12).
We first converted the spike-sorted units to estimated firing rate matrices by convolving the spike trains with a 30 ms Gaussian kernel (Fig. 7). Applying CrEIMBO to the firing rate data with a maximum of $K=12$ sub-circuits and $p_{j}=7$ ensembles (Fig. 13) per region, we found multi-regional motifs with variability across sessions.
@@ 4.2 CrEIMBO Discovers Multi-Regional Dynamics in Human Electrophysiology Data
We then tested CrEIMBO on the first $18$ seconds from whole sessions together with a maximum of $K=12$ sparse sub-circuits and as most $p_{j}=7$ ensembles per region (full parameter list available in Tab. 2). First we note that CrEIMBO can accurately reconstruct the data with high accuracy across all sessions (Fig. 3A, B). Interestingly, the ensemble compositions identified by CrEIMBO (Fig. 3E) presented sparse patterns with one dense ensemble for most regions that we hypothesize captures the ‘general’ mean field activity of that region (the right-most column of each diagonal block in Fig. 3E). The sparse ensembles we hypothesize capture the more specialized, nuanced functionality. The ensemble dynamics ($\bm{X}^{d}$, Fig. 3G) reveal that some ensembles are constantly active, implying that they might play an important role in the observed neural activity.
@@
The identified set of global sub-circuits ($\{\bm{f}_{k}\}_{k=1}^{K}$) exhibits distinct localized motifs (Fig. 3H), i.e., most sub-circuits present clear trends of either sourcing or targeting the same area. This is in contrast to other approaches (Sec. F), that yield overlapping sub-circuits (Fig. 19) that emphasize only limited circuitry (Fig. 20). Interestingly, sub-circuits 10 and 12 ended up almost empty (mainly zero-ish values, Fig. 3H, 16). This highlights CrEIMBO’s ability to automatically nullify extra or redundant sub-circuits—i.e., a form of model selection—via the sparsity regularization. The number of active connections and average connection strength (Fig. 7J, K) present modest variability between sub-circuits. This underscores the model’s ability to identify distinct sub-circuits with different structural behaviors, capturing various interaction resolutions. Specifically, it can find sub-circuits of varying strengths and is not limited to only k-sparse sub-circuits. From a neuroscience perspective, this suggests that underlying sub-circuits can vary in density and the number of participating areas and ensembles.
Interestingly, most identified interactions occur between distinct areas (i.e., inter-region interactions) with either multiple ensembles from the same area affecting together ensembles from a different area (e.g., $\bm{f}_{3}$), or multiple ensembles from diverse areas converging to the same target area (e.g., $\bm{f}_{5}$). However, these source or target groups of ensembles vary across sub-circuits, hence the full repertoire of activity cannot be captured by a single. In addition to these between-area interactions, we also observed, although to a lesser degree, within-area interactions (e.g., as in $\bm{f}_{8}$ and $\bm{f}_{11}$, as marked by black arrows). These dynamics emphasize that within-area amplification or regulation dynamics do exist in addition to global cross-region computations, however seem to be less salient than cross-area communication.
When exploring the activation patterns of these sub-circuits (Fig. 3D, Fig. 17), we reveal variability between subjects and sessions. However, the usage of overall similar sub-circuits, with some sub-circuits not being used at all in any of the exemplary sessions (Fig. 3C, F, e.g., $\bm{f}_{12}$ not being used), while others are constantly used with high variability over time (e.g., $\bm{f}_{8}$ light-green). Interestingly, when exploring the within-session pairwise correlations between sub-circuits coefficients (Fig. 7I), it seems that most coefficients present very low ($<0.1$) correlations, indicating that the activity of the corresponding circuits is not similar and they might capture different cognitive processes, while other pairs of sub-circuits may show similarities in activity. This might indicate that for a specific subject or session, they may be related. For instance, a specific image may evoke some emotional response in one subject but not in another (e.g., $\bm{c}_{9}$ $\bm{c}_{11}$ in the second session of subject 10, however not in the first session of that subject nor in the average correlations across all sessions).
5 Discussion.
In this paper we introduced CrEIMBO (Cross Ensemble Interactions in Multi-view Brain Observations), a novel approach for uncovering multi-regional dynamics in neural data collected across multiple sessions. CrEIMBO addresses the challenge of integrating non-simultaneous neural recordings by leveraging joint inference and sparsity regularization to capture the underlying neural sub-circuits governing brain activity.

### v2

@@ 3 Problem Introduction and Approach
Fitting this model requires identifying the global operators $\{\bm{f}_{k}\}_{k=1}^{K}$ and ensembles $\{\bm{A}^{d}\}_{d=1}^{D}$. For this, we employ an alternating approach and iterate between updating $\{\bm{f}_{k}\}_{k=1}^{K}$ and inferring, for each session $d$, 1) the ensemble compositions $\bm{a}:=\bm{A}^{d}$, 2) their dynamics $\bm{x}$, and 3) the sub-circuits’ temporal coefficients $\bm{c}_{t}$. These steps are iterated until convergance.
@@
Ensemble update: We update $\bm{a}$ per-row (neuron) $n$ ($\bm{a}_{n}$). Following the work of Mudrik et al. (2024b), we infer the sparse ensemble structures with graph-driven re-weighted $\ell_{1}$ regularization that groups together neurons with shared activity patterns, while pushing apart neurons who do not behave similarly. The graph for each area $j$, $\bm{h}^{j}\in\mathbb{R}^{n_{j}\times n_{j}}$, is calculated based on a data-driven Gaussian kernel that measures the temporal similarity between neurons, such that $\bm{h}^{j}_{n1,n2}=exp(\frac{\|\bm{y}^{j}_{n1}-\bm{y}^{j}_{n2}\|_{2}^{2}}{\sigma_{h}})$ where $\sigma_{h}$ is a hyperparameter. Specifically, for each row $\bm{a}_{n,:}$ and considering only the ensembles in region $j_{n}$ that neuron $n$ is part of,
$\widehat{\bm{a}}_{n,:}=\arg\min_{\bm{a}_{n:}}\|\bm{y}_{n:}-\bm{a}_{n:}\bm{x}_{j_{n}:}\|_{2}^{2}+\sum_{j=1}^{p}\bm{\lambda}^{d}_{n,j}|\bm{a}_{n,j}|\textrm{ where }\bm{\lambda}_{n,j}=\frac{\beta_{1}}{\beta_{2}+|\widehat{\bm{a}}_{n,j}|+\beta_{3}|\bm{h}_{n:}^{j}\widehat{\bm{a}}_{:j}|}.\vspace{-5pt}$ (1)
@@
Here, $\bm{x}_{j_{n},:}$ is the activity of the ensembles of the $j_{n}$-th region and $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are scalars that control the effect of the graph on the regularization. The graph adjacency $\bm{h}_{n:}\widehat{\bm{a}}_{:j}$ weights the sparse membership in the $j$-th ensemble based on temporal similarity. Specifically, if $\bm{h}_{n:}\widehat{\bm{a}}_{:j}$ is high (i.e., the neighbors neuron $n$ on the graph are in the current estimate of the $j$-th ensemble), $\bm{\lambda}_{n,j}$ is small, resulting in lower sparsity regularization on neuron $n$ for that ensemble, promoting it to be included. Alternatively, if this correlation is small, the regularization weight $\bm{\lambda}_{n,j}$ is large, and neuron $n$ will be more likely not to be included in ensemble $j$.
@@
${\widehat{\bm{x}}_{t},\widehat{\bm{c}}_{t}=\arg\min_{\bm{x}_{t},\bm{c}_{t}}\|\bm{y}_{t}-\bm{a}\bm{x}_{t}\|_{2}^{2}+\left\|\bm{x}_{t}-\left(\sum_{k=1}^{K}\bm{f}_{k}\bm{c}_{k,(t-1)}\right)\bm{x}_{t-1}\right\|_{2}^{2}+\lambda_{c}\|\bm{c}_{t}\|_{1}}$ (2)
where $\lambda_{c}\in\mathbb{R}$ is sparsity-regularization weights on the basis interactions’ activations.
@@
$\bm{F}^{\textrm{all}}=\arg\min_{\bm{F}^{\textrm{all}}}\|\bm{x}^{+}-\bm{F}^{\textrm{all}}(\bm{cx})\|_{2}^{2}+\lambda_{F}\|\text{vec}(\bm{F}^{\textrm{all}})\|_{1}+\lambda_{\rho}\sum_{{k_{1}},{k_{2}},({k_{1}}\neq{k_{2}})}\rho({\bm{f}}_{k_{1}},{\bm{f}}_{k_{2}})$ (3)
@@
where $\bm{F}^{\textrm{all}}\in\mathbb{R}^{p\times pK}$ is an horizontal concatenation of all $\bm{f}$s, and $\bm{x}^{+}\in\mathbb{R}^{p\times\sum_{d\in\textrm{batch}\mathrm{T}^{d}}}$ is the horizontal concatenation of all $\{\{\bm{x}_{t+1}^{d}\}_{t=1}^{\mathrm{T}^{d}}\}_{d\in\text{batch}}$, where “batch” means that this operation is in practice taken on a random subset of sessions for computational complexity considerations. Above, $\bm{cx}_{t}^{\textrm{batch}}\in\mathbb{R}^{Kp\times 1}=[(\bm{c}^{d}_{t}\otimes[1]_{1\times p})\circ([1]_{1\times K}\otimes\bm{x}^{d}_{t})^{T}]^{T}$, such that $\bm{cx}\in\mathbb{R}^{Kp\times\sum_{d\in\text{batch}}\mathrm{T}^{d}}$ is the horizontal concatenation of all $\{\{(\bm{cx})_{t}^{d}\}_{t=1}^{\mathrm{T}^{d}}\}_{d\in\text{batch}}$. The operator $\text{vec}(\cdot)$ flattens a matrix to a vector and $\lambda_{F}$ is $\ell_{1}$ sparsity-promoting weight on $\bm{F}^{\textrm{all}}$. $\rho(\cdot)$ represents correlation, such that $\sum_{{k_{1}},{k_{2}}}\rho(\bm{f}_{k_{1}},\bm{f}_{k_{2}})$ is a de-correlation term with weight $\lambda_{\rho}$ used to ensure that distinct $\bm{f}$s are not too similar. See Algorithm 1 for a method summary and Section E for complexity analysis.
Figure 2: A-E: Testing CREIMBO over 312 repeats with varying random initializations and random seeds (Sec. B) reveals high correlations between ground truth and fit components for all unknowns, $\bm{Y}$, $\bm{A}$, $\bm{x}$, $\bm{F}$, $\bm{c}$ (subplots A-E respectively). F-G:Ablation Experiment: Comparison of the ensemble trajectory ($\bm{x}^{d}$, subplot F) and the ensemble compositions ($\bm{A}$, subplot G) for the $1$-st session as identified by CREIMBO vs. various ablations (details in Sec. D, see Figure 7 for more sessions). H-J: Comparing CREIMBO to ablation variants in terms of $-\log(\text{MSE})$ between ground truth and identified dynamics, ensembles, and sub-circuits. “Single Session” refers to training CREIMBO on sessions individually. Details in Sec. D. K: Baseline Comparisons: the latent dynamics identified by the baselines (SDLS, rSLDS (Linderman et al., 2016), multi-population rSLDS (Glaser et al., 2020), details in Sec. D) for the first two sessions (merged sessions baselines provide a single trajectory). L-O: Comparison of the latent dynamics and identified ensemble matrices compared to the ground truth across the baselines.
@@
Figure 3: CREIMBO identifies cross-regional neural sub-circuits underlying multi-session human brain recordings. A: Two exemplary observations ($\bm{Y}$, right) compared to their reconstruction by CREIMBO ($\widehat{\bm{Y}}$, left). B: Reconstruction performance in terms of $R^{2}$ and relative error across all sessions. C: The sub-circuit coefficients for 3 exemplary subjects. D: The standard deviation (std) of the coefficients over time across the two exemplary sessions. E: Exemplary identified sparse ensemble matrices ($\{\bm{A}^{d}\}_{d=1}^{D}$) for the observations from A. F: Relative std of the coefficients across four subjects reveals usage of similar sub-circuits. G: The identified ensembles’ activity ($\bm{X}^{d}$) for the two exemplary sessions. H: The identified sub-circuits. Width indicates effect magnitude. Each circle corresponds to an ensemble with its color indicating the ensemble’s area. Black arrows near $\bm{f}_{8}$ and $\bm{f}_{11}$ highlight within-region interactions. I: Pairwise correlations between within-session sub-circuit coefficients (left, middle) and average within-session correlations across all sessions. STD marked by “+” marker size. J: Mean and STD of connections $f$s. K: Number of non-zero connections per $\bm{f}_{k}$.
@@ 4 Experiments
CREIMBO recovers Ground Truth Components in Synthetic Data: We first assess CREIMBO’s ability to recover ground-truth components from synthetic data. We generated $K=3$ sub-circuits represented by a set of 3 rotational matrices (Fig. 5B). These sub-circuits capture interactions among $p=3$ ensembles from $J=3$ regions ($p_{j}=1$ for all $j=1,\dots,J$). Each region consists of a random number of neurons drawn from a uniform distribution between $4$ to $9$. We finally generated the data by simulating 5 sessions with different neuron identities and varying numbers of neurons per region(Fig. 5A and D). CREIMBO was able to recover the ground truth components with high accuracy, as measured by correlation with the ground truth, across $312$ random noisy initialization repeats with random parameters from a range described in 1. Accurate recover included all parameters: the reconstructed observations $\bm{y}$ (Fig. 2A), ensemble composition $\bm{A}$ (Fig. 2B) and activity $\bm{x}$ (Fig. 2C), sub-circuits $\bm{F}$ (Fig. 2D), and coefficients $\bm{c}$ (Fig. 2E). Comparing CREIMBO’s results to components identified under diverse ablations (details in D), we found that approaches that either separate the ensembles and dynamics identifications steps, analyze each session separately, or do not consider the localized structure of the ensembles, fail to recover the ground-truth components, i.e., the resulting fit model parameters are less correlated with the ground-truth observations (Fig. 2F-J, 7). We further evaluated CREIMBO’s ability to recover hidden components compared to other methods, (Fig. 2 K-O, Fig.8. baselines details in Sec. D). While the latter shows improved performance, CREIMBO more accurately recovers the ground truth hidden dynamics and ensembles. Importantly, these models are not designed to identify multiple co-occurring processes or sparse ensembles, in contrast to CREIMBO, and they are not suitable for multi-sessions with varying neural identities, which limits the comparison from the outset.
@@
Next, we tested CREIMBO on a richer synthetic example with $D=15$ sessions. Each session had a varying number of neurons (a random integer between $14$ and $19$ per region from a uniform distribution) and distinct sub-circuit coefficients (see Fig. 11 for the generated ground truth components and 9 for regional distributions and masking). The sub-circuits were set as rotational matrices, similar to the previous experiment. Over $204$ random initializations with different parameters (see Sec. B) we again found that CREIMBO robustly identifies the ground truth latent dynamics, ensemble compositions, and sub-circuits. Moreover, CREIMBO accurately reconstructed the observations with high correlation to the ground truth across all initializations (Fig. 11E) with low relative error (Fig. 11F). We further tested CREIMBO on more advanced data that included $D=40$ sessions with a maximum of $4$ regions and $p_{j}=3$ ensembles per region (Fig. 12). We found that CREIMBO manages to recover the components with high correlation with the ground-truth (Fig 13).
@@
CREIMBO Discovers Multi-Regional Dynamics in Electrophysiology Data: We then tested CREIMBO on human neural recordings from a high-density electrode array provided by Kyzar et al. (2024). The data consists of neural activity from overall $p=10$ brain areas with limited cross-session regional overlap (Fig. 14, 15), recorded while subjects performed a screening task (details in Kyzar et al. (2024)). The data encompasses $21$ subjects across $D=41$ non-simultaneous sessions. For each subject (except Subject 19), the data offers 1) a “Screening” session, and 2) a “Sternberg test” session (details in Kyzar et al. (2024)). We first converted the spike-sorted units to firing rate matrices by convolving the spike trains with a $30$ms Gaussian kernel (Fig. 10), and then tested CREIMBO on all sessions together with a maximum of $K=12$ sub-circuits and as most $p_{j}=7$ ensembles per region (full parameter list in Tab. 2). First we note that CREIMBO can accurately reconstruct the data with high accuracy across all sessions (Fig. 3A, B). Interestingly, the ensemble compositions identified by CREIMBO (Fig. 3E) present sparse patterns with one dense ensemble for most regions that we hypothesize capture the “background” mean field activity of that region (right-most column of diagonal blocks in Fig. 3E). The sparse ensembles we hypothesize capture the more specialized, nuanced functionality. The ensembles activity ($\bm{X}^{d}$, Fig. 3G) reveal some constantly-active ensembles, implying these neuron groups are important for neural processing.
@@
The universal sub-circuits ($\{\bm{f}_{k}\}_{k=1}^{K}$) exhibit distinct localized motifs (Fig. 3H), i.e., most sub-circuits present clear trends of either sourcing or targeting the same area. This is in contrast to other approaches (Sec. D), that yield overlapping uninterpretable sub-circuits (Fig. 22) that emphasize only limited circuitry (Fig. 24). Interestingly, sub-circuits $10$, $12$ ended up almost empty (mainly zero-ish values, Fig. 3H, 19). This highlights CREIMBO’s ability to automatically nullify redundant sub-circuits—i.e., a form of model selection—via the sparsity regularization. The number of active connections and average connection strength (Fig. 10J, K) show modest variability between sub-circuits, highlighting the model’s ability to identify distinct sub-circuits with varying interaction resolutions, including those of varying strengths rather than being limited to just k-sparse sub-circuits. From a neuroscience perspective, this suggests that underlying neural sub-circuits can vary in density and the number of participating areas and ensembles. Interestingly, most identified interactions occur between distinct areas (i.e., inter-regional interactions) with either multiple ensembles from the same area affecting together ensembles from a different area (e.g., $\bm{f}_{3}$), or multiple ensembles from diverse areas converging to the same target area (e.g., $\bm{f}_{5}$). However, these source or target groups of ensembles vary across sub-circuits, hence the full repertoire of activity cannot be captured by a single one, highlighting the importance of CREIMBO’s capacity to disentangle the combined activity of multiple circuits. We also observe, but to a lesser degree, within-area interactions (e.g., black arrows in $\bm{f}_{8}$, $\bm{f}_{11}$). These dynamics emphasize that within-area amplification or regulation do exist in addition to global cross-region computations. When exploring the activation patterns of these identified sub-circuits (Fig. 3D, Fig. 20), CREIMBO reveals cross-subject and cross-session variability. Notably, some sub-circuits are consistently used with high variability over time (e.g., $\bm{f}_{8}$ light-green), while others, such as $\bm{f}_{12}$, are not utilized at all in any exemplary session (Fig. 3C, F, e.g., $\bm{f}_{12}$). Moreover, when exploring sub-circuits with high activity variability across sessions (Fig. 3D,F), we observe that some are used within-subjects but across both tasks (e.g., $\bm{f}_{2}$ in subject 14, Fig. 3D), while others are shared across both subjects and tasks (e.g., $\bm{f}_{11}$, used by subjects 10 and 14 in both task phases). The analysis of within-session pairwise correlations between sub-circuit coefficients (Fig. 10I) reveals low correlations ($<$0.1) for most pairs. This indicates that the activities of distinct sub-circuits differ and reflect different cognitive processes.
Since ground-truth components are unavailable in real data, direct validation of such components is impossible. Hence, we sought a proxy to evaluate CREIMBO’s performance on real-data by testing its consistency under increasing observation noise levels. Particularly, we re-ran CREIMBO on the human recordings after adding increasing variance i.i.d Gaussian noise (with $K=8,p_{j}=6$ Fig.23). CREIMBO remains robust to increasing noise, and experiences a phase transition at a specific noise level ($\frac{\sigma_{\text{noise}}}{\sigma_{\text{data}}}\sim 0.2$,Fig.23 A,B, E, F,H,I), which align with the dictionary-learning literature (Studer & Baraniuk, 2012). In this noise experiment, CREIMBO further identifies that within-area ensemble composition correlations (Fig.23C,D) remain stable until a noise level of around $\sigma_{\text{noise}}\approx 10$, after which these correlations weaken as noise increases, which may imply that high noise in data can obscure meaningful relationships between ensembles. Additionally, it identified that the dynamics’ coefficients remain robust until $\sigma_{\text{noise}}\approx 1.8$, while exhibiting an increasing internal frequency as the noise increases, potentially to account for noise that does not follow dynamical rules (Fig. 23G,  25). This suggests that frequency of fast transitions in CREIMBO’s ${\bm{c}_{k,t}}$ could indicate noise in the system that does not adhere to the underlying dynamics.
CREIMBO Discovers Regional Interactions Predictive of Task Variables: We tested CREIMBO’s ability to infer task-related variables from mice whole-brain Neuropixels multi-session data during a memory-guided movement (data from Chen et al. (2024), Fig. 2627). CREIMBO identifies intra- and inter-area brain interactions via the sub-circuits (Sec. F, Fig. 29C,D), including cross-regional flows into or from key areas responsible for memory (from $\bm{f}_{5}$ and to $\bm{f}_{8}$ the hippocampus), planning (from frontal cortex, e.g. $\bm{f}_{3}$), and movement (from primary motor cortex, $\bm{f}_{4}$), which are needed for the task. Within-area sub-circuits (Fig. 29D) further showed within-area ensemble interactions (e.g., the basal ganglia in $\bm{f}_{1}$ and the secondary motor cortex in $\bm{f}_{2},\bm{f}_{7}$, along with self-activation/inhibition of ensembles in other sub-circuits (e.g. $\bm{f}_{4}$). Moreover, CREIMBO’s dynamic coefficients (${\bm{c}_{kt}}$) captured task-related patterns across trials and sessions (Fig. 4A,B). When training a regularized logistic-regression model, we found that we could predict from CREIMBO’s circuits activations ($\bm{c}_{kt}$) various task variables including outcome, early lick, and lick side far above chance levels (Fig. 4C-J, p-values < $1\times 10^{-10}$). Furthermore, CREIMBO’s sub-circuit activations across different task periods highlight how specific multi-regional interactions capture different aspects of the task. For instance, towards the trial end ($t_{3}$ window), which includes the lick movement, $c_{7}$ shows increased importance (Fig. 4E). Notably, $c_{7}$ corresponds to the activity of the $\bm{f}_{7}$ sub-circuit that captures flows to the secondary motor cortex (Fig. 29C). Another example is $c_{8}$ (activity of sub-circuit $\bm{f}_{8}$ that includes flows into hippocampus 29C), which shows increased feature importance in the first time window $t_{0}$. This aligns with the potential involvement of memory-related processes required with the stimulus appearance. Thus CREIMBO demonstrates the ability to capture interpretable neural dynamics and predict task variables that reveal complex regional interactions. (see Sec. F for more detailed results).
Figure 4: Results of task-variable prediction using CREIMBO’s dynamic coefficients as input. Prediction was done based on coefficients from all used trials and sessions (19 sessions, 40-60 trials each, see Sec. F). A: Dynamic coefficients from example traces, shade marks within trial, bordered by starting and ending point. B: Dynamic coefficients for correct (“hit”) vs. wrong (“miss”) trials. C: Accuracy score of predicting task variables based on coefficients vs. chance. D, F,G,H,I: Confusion matrices of predicting varying task variables based on dynamic coefficients. E: importance of the different coefficients and time points for prediction.
5 Discussion, Limitations, and Future Work